change 74(in base 10)to base 4.
Answer by stanbon(75887)

change 74(in base 10)to base 4.
Successive divisions of 74 by 4 will give
you remainders that are:
1st the units digit
2nd the 4's digit
3rd the 4^2 digit
4th the 4^3 digit
etc.
------------------
74/4 = 18 with R = 2
18/4 = 4 with R = 2
4/4 = 1 with R = 0
---
74(base 10) = 1022(base 4)
=================================
Cheers,
Stan H.